OpenStudy (anonymous):

Students in an English class took a final exam. They took equivalent forms of the exam at monthly intervals thereafter. The average score S(t), in percant, after t months was found to be given by S(t)=68-20log(t+1),t greater than or equal to 0. a) what was the average score after 4 months? after 24 months? Round your answers to a whole percent, if necessary.

OpenStudy (zarkon):

just evaluate your function at t=4 and t=24

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The formula for calculating the amount of money returned for an initial deposit into a bank account or CD(certificate of deposit) is given by: A=P(1+r/n)^nt: A is the amount of the return. P is the principal amount initially deposited. r is the annual interest rate (expressed as a decimal). n is the number of compound periods in one year. t is the number of years. Carry all calculations to six decimal places on each intermediate step, then round the final answer to the nearest whole cent. Suppose you deposit $2,000 for 5 years at the rate of 8%. a) Calculate the return (A) if the bank compounds annually (n=1). Round your answer to the nearest whole cent. Answer: b) Calculate the return(A) if the bank ocmpounds quarterly (n=4). Round your answer to the nearest whole cent. c) if a bank compounds continously, then the formula used is A=Pe^rt where e is a constnant and equals approximately 2.718282. Calculate A with continuous compunding. Round your answer to the nearest whole cent. PLEASE HELP!!!

OpenStudy (zarkon):

\[A=P\left(1+\frac{r}{n}\right)^{nt}\] \[A=2000\left(1+\frac{.08}{1}\right)^{1\times5}\]

OpenStudy (anonymous):

how do I solve with the exponent do I simplify 1*5 and then solve in the parenthesis and then find the answer?

OpenStudy (zarkon):

\[A=2000\left(1+\frac{.08}{1}\right)^{1\times5}=2000(1.08)^5\]

OpenStudy (anonymous):

:) You're nice.. thanks.. I really needed the help. :)

OpenStudy (zarkon):

for (b) \[A=2000\left(1+\frac{.08}{4}\right)^{4\times5}\]

OpenStudy (anonymous):

so that's 2000(1.02)^20 right?

OpenStudy (anonymous):

which rounded to the nearest whole dollar would make it $2971.89

OpenStudy (anonymous):

i mean whole cent.. not whole dollar..

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

now part c.. that's not like the other two.. A=2000(2.718282)^ dunno how to get r and t would be 5.. i think.. just a little help with that.. then i'm good.. i think.. :)

OpenStudy (zarkon):

when you let \[n\to \infty\] then \[A=P\left(1+\frac{r}{n}\right)^{nt}\] transforms into \[A=Pe^{rt}\]

OpenStudy (zarkon):

then we have \[A=2000e^{.08\times5}\]

OpenStudy (anonymous):

so.. e is that 2.718282 number? multiplied by 2000?

OpenStudy (zarkon):

yes \[e\approx 2.718281828\]

OpenStudy (anonymous):

ok.. i know how to do it now.. Thank God for you!!! :)